OpenStudy (anonymous):

Solve the following logarithmic equations a) log (2x-3) – log 3=0 b) ln(8x-4) = ln 2+lnx

OpenStudy (experimentx):

use the property of log \[ \log a + \log b = \log ab \] \[ \log a - \log b = \log \frac{a}{b} \] \[ \log 1 = 0\]

OpenStudy (anonymous):

what is about the answer of the second one?

OpenStudy (experimentx):

x = 2/3